The square of 9 less than a number is 3 less than the number. What is the number?

Answer by ikleyn(52776)

.
The square of 9 less than a number is 3 less than the number. What is the number?
~~~~~~~~~~~~~~~~~~~

Let n denotes a number under the problem's question.


Equation

    (n-9)^2 = n-3


Simplify and find n

    n^2 - 18n + 81 = n-3

    n^2 - 19n + 84 = 0

    (n-12)*(n-7) = 0


The roots are  n= 7  and n= 12.


Both the roots satisfy the problem's description, so they both are the solutions to the problem.

Solved.

Actually, the problem's description is ambiguous and can be read and interpreted in different way;
I presented here one of possible interpretations with the solution.